A and B contribute $140000 and $180000 respecrively in a business partnership. Of the profit, A receives 20% as manager and the rest is shared in the ratio of their capitals. What is the ratio of A's total share of the profit to B's?
Answer by Theo(13342)

A contributes 140,000
B contributes 180,000
total contribution is 320,000.

A proportion of total contribution = 140 / 320 = .4375
B proportion of total contribution = 180 / 320 = .5625

A receives 20% of the profit as manager.

the rest is shared in the same proportion as their contributions.

if the profit is x, then:

A gets .2 * x + .4375 * .8 * x = .55 * x
B gets .5625 * .8 * x = .45 * x

for example:

assume the profit is 400,000.

A gets .20 * 400,000 = 80,000 + .4375 * .8 * 400,000 = 140,000 for a total of 140,000 + 80,000 = 220,000.

B gets .5625 * .8 * 400,000 = 180,000

the total profit is 400,000
A gets 220,000 = 220/400 = .55 * the total profit.
B gets 180,000 = 180/400 = .45 * thee total profit.

the ratio of A's total share of the profit to B's total share of the profit is equal to .55 / .45 = 11/9.

if the total profit is broken up into 20 parts, A gets 11 of them and B gets 9.

11/20 = .55
9/20 = .45

i think that's what you're looking for.

the ratio of A's total share of the profit to B's total share of the profit is 11/9.